# Your Latent Reasoning is Secretly Policy Improvement Operator

### v3

@@ 1 Introduction
The goal of reasoning models is to start from a set of specific examples or observations and infer a general rule or pattern. Such models systematically manipulate and connect pieces of information to infer new conclusions and solve problems not explicitly stated in the training data.
Recently, the use of latent reasoning and iterative refinement loops has become a major driver of progress in deep learning reasoning. Looped transformers (Giannou et al., 2023; Yang et al., 2023) repeatedly apply the same transformer block with input injection at each step and achieve better performance than a standard one forward pass transformer on reasoning and meta learning tasks, why using 10x less number of parameters (Yang et al., 2023).
@@
TRMs are difficult to train efficiently: supervision is primarily applied to the final output, leaving early refinement steps with weak credit assignment and making optimization brittle in the small-data regime. We propose Deep Improvement Supervision (DIS), a training method that supplies step-wise intermediate targets during recursion. In our main implementation, these targets are generated by a monotonic discrete corruption schedule of the ground-truth output, turning each recursion step into a supervised subgoal.
@@ 2.1 Hierarchical Reasoning Models
$\mathbf{z}^{n+1}_{L}$ $\leftarrow f^{\phi}_{L}((\mathbf{z}^{n}_{L}+\mathbf{x})+\mathbf{z}^{n}_{H}),\quad\text{(repeated $n$ times)}$ (1)
$\mathbf{z}^{n+1}_{H}$ $\leftarrow f^{\psi}_{H}(\mathbf{z}^{n+1}_{L}+\mathbf{z}^{n}_{H}),$
@@ 2.2 Tiny Recursive Models
The TRM paper explicitly reframes this: $\mathbf{z}_{H}$ is simply the current (embedded) solution… $\mathbf{z}_{L}$ is a latent feature that does not directly correspond to a solution but can be transformed into one by $f_{H}$”. It was shown on the Sudoku example that when you reverse-embed + argmax, the tokenized $\mathbf{z}_{H}$ looks like the solved grid, while tokenized $\mathbf{z}_{L}$ is not realted to the solution.
@@ 2.3 Policy Improvement Algorithm
Recently was shown that policy improvement framework does not necessarily rely on explicitly learning a value function. Recent work formalizes a tight connection between classifier-free guidance (CFG) in diffusion/flow models and policy improvement in reinforcement learning (Frans et al., 2025). This work provides a framework for the analysis of diffusion models as RL methods. For a state $\mathbf{s}$ and an action/output $\mathbf{a}$, the classifier-free guidance RL (CFGRL) method parameterizes a target policy as:
@@ 3 Formalization of Latent Reasoning as Policy Improvement
Both (TRMs) and (HRMs) are often motivated by effective depth: reusing a small parameter-shared network is said to emulate very deep computation with few parameters. In HRMs this story is tied to Deep Equilibrium (DEQ) models, where one assumes convergence to a fixed point and invokes the Implicit Function Theorem to justify differentiating through only the last iterate (Bai et al., 2019). However, follow-up analyses indicate that the fixed-point assumption is typically violated in practice: latent residuals remain non-negligible under truncation, and a one-step DEQ approximation does not fully explain the empirical gains (Jolicoeur-Martineau, 2025).
@@
$s_{t}\;=\;(x,z_{L}^{t},z_{H}^{t}),$
where $x$ is the embedded input, $z_{L}^{t}$ is a latent “scratchpad”, and $z_{H}^{t}$ is the current solution embedding. A parameter-shared backbone is used in two roles:
$z_{L}^{t+1}=f_{\phi}^{(L)}(x,z_{H}^{t},z_{L}^{t}),\qquad z_{H}^{t+1}=f_{\phi}^{(H)}(z_{H}^{t},z_{L}^{t+1}),$ (6)
and an output head $f_{O}$ maps $z_{H}$ to logits over the next token/action set $\mathcal{A}$. Only $z_{H}$ is directly decoded and trained with cross-entropy.
@@ 3.1 Two policies per recursion answer
Reference policy (pre-reasoning). Decode before updating:
$\ell_{u}^{t}\;:=\;f_{O}(z_{H}^{t}),\qquad\hat{\pi}_{t}(a\mid s_{t})\;:=\;\mathrm{softmax}(\ell_{u}^{t})[a].$ (7)
@@
$\ell_{c}^{t}\;:=\;f_{O}(z_{H}^{t+1}),\qquad\pi_{t}^{+}(a\mid s_{t})\;:=\;\mathrm{softmax}(\ell_{c}^{t})[a].$ (8)
Intuitively, $\hat{\pi}_{t}$ is the model’s first guess at step $t$, while $\pi_{t}^{+}$ is the refined distribution after one reasoning step.
We now introduce the same object used in policy-improvement analyses: an optimality (or improvement) variable. Let $o\in\{0,1\}$ indicate whether an action $a$ is “good” for state $s$. (For supervised next-token prediction, the simplest choice is $o=1$ iff $a$ equals the ground-truth token; more general choices can encode task-level reward or self-consistency.) A standard KL-regularized improvement step produces a new policy by reweighting a reference policy $\hat{\pi}$ using an optimality likelihood $p(o=1\mid s,a)$:
@@
is an improvement over $\hat{\pi}$. Please, see Remark 1 and Theorem 1 (Frans et al., 2025). We can say that, (11) is a special case of (12) with $f(u)=\exp(wu)$.
@@ 3.2 Improvement from a conditional policy
i.e., “the policy after one reasoning step behaves like the policy conditioned on the event that the answer is improved.”
Under this interpretation we can recover an implicit optimality model by Bayes’ rule:
@@
Equation (16) is the policy-improvement analogue of the “Bayes inversion” step: it rewrites a multiplicative optimality reweighting in terms of a pair of policies (unconditional vs. optimality-conditioned). Concretely, latent reasoning provides both factors $\hat{\pi}_{t}$ and $\pi_{t}^{+}$, and therefore implicitly defines the entire one-step improvement family $\{\pi_{t,w}\}_{w\geq 0}$.
From (14), the improvement score is the log-ratio
@@ 3.3 A supervised improvement criterion
For next-token prediction, let $y^{\star}$ be the ground-truth token at the current position. Define the cross-entropy under the improved policy family (16):
@@
Thus increasing $w$ (trusting the reasoning step more strongly) decreases the loss if and only if the ground-truth action has above-average improvement score:
@@ 4 Deep Improvement Supervision
Section 3 reframes a latent-reasoning step as a policy-improvement operator: each recursion produces a reference policy $\hat{\pi}_{t}(\cdot\mid s_{t})$ and a post-reasoning policy $\pi_{t}^{+}(\cdot\mid s_{t})$, and their log-ratio $A_{t}(s_{t},a)=\log\pi_{t}^{+}(a\mid s_{t})-\log\hat{\pi}_{t}(a\mid s_{t})$ acts as an advantage-like improvement signal. The analysis yielded a concrete Advantage Margin condition: a latent update helps exactly when the ground-truth action has above-average improvement score under the interpolated policy family $\pi_{t,w}$ (Eq. (18)).
@@
In a recursive architecture, the input to step $s$ (which produces $\ell_{u}^{s}$) is a deterministic function of the post-reasoning state of the previous step (which produces $\ell_{c}^{s-1}$). Under teacher-forced recursion, training $\ell_{c}^{s-1}$ toward $\mathbf{y}^{\dagger}_{s-1}$ implicitly anchors $\ell_{u}^{s}$ near the previous target. Thus we use the single-loss
@@ 4.1 DIS as direct optimization of the Advantage Margin
Training $\ell_{c}$ on $\mathbf{y}^{\dagger}_{s}$ while $\ell_{u}$ is anchored to $\mathbf{y}^{\dagger}_{s-1}$ forces the residual $\Delta\ell^{s}=\ell_{c}^{s}-\ell_{u}^{s}$ to encode the incremental improvement direction from $\mathbf{y}^{\dagger}_{s-1}$ to $\mathbf{y}^{\dagger}_{s}$. Equivalently, in probability space, DIS shapes the log-ratio advantage $A_{s}(a)=\log\pi_{s}^{+}(a\mid s)-\log\hat{\pi}_{s}(a\mid s)$ assign higher-than-average advantage to the next target token(s), aligning with the theoretical Advantage Margin condition (Eq. (18)). This converts the previous section’s existence/diagnostic statement (“improvement occurs iff the margin is positive”) into a constructive training rule: make the margin positive at every step by supervising the post-reasoning policy on a strictly improving target sequence.
@@ 4.2 Algorithm
Standard TRM training backpropagates a single terminal supervision signal through the entire recursive chain, so the model must simultaneously (i) discover useful intermediate states and (ii) align each latent update with improvement. DIS isolates the second requirement by supervising the improvement operator itself: each recursion step is trained to map “previous target” $\mathbf{y}^{\dagger}_{s-1}$ to “next target” $\mathbf{y}^{\dagger}_{s}$, directly instantiating the stepwise improvement interpretation from Section 3.
@@
    Programmable edits. A deterministic generator creates a path from input to output, e.g., puzzle solvers that reveal a constraint-consistent move per step, yielding $\mathbf{y}^{\dagger}_{s+1}=\mathrm{Edit}(\mathbf{y}^{\dagger}_{s})$.
@@
    Discrete corruption schedule. Define a corruption process $q_{\beta}(\tilde{\mathbf{y}}\,|\,\mathbf{y}^{\star})$ (e.g., token masking or random replacement with rate $\beta$). Choose a decreasing noise schedule and sample $\mathbf{y}^{\dagger}_{s}\sim q_{\beta_{s}}(\cdot\,|\,\mathbf{y}^{\star})$ so that the targets become progressively less corrupted, yielding a simple monotone-improvement curriculum. [^1] We do not position DIS as diffusion modeling: DIS does not require diffusion schedulers, score matching, or explicit density-ratio estimation. A corruption schedule is merely a convenient way to generate monotone intermediate targets in discrete spaces (Lou et al., 2023).
@@
During training, the model runs the TRM recursion for $N_{\text{sup}}$ supervision steps. At each step $s$, it computes the post-reasoning logits $\ell_{c}^{s}$ and applies $\mathrm{CE}(\ell_{c}^{s},\mathbf{y}^{\dagger}_{s})$. By the theory in Section 3, this stepwise supervision trains each latent update to increase the relative log-probability of the next improved target – i.e., to implement a non-trivial policy improvement operator rather than an unconstrained recurrent computation.
@@ 5 Experiments
From algorithmic perspective, we incorporate the time step $t$ into the model input and observe that an integer-based time index $(0,1\dots,N_{sup})$ yields superior results compared to standard continuous diffusion time conditioning $t\in[0,1])$. The total simplification that we made in comparison to the original TRM are:
@@
  • Halting/ACT: Although TRM/HRM use halting where HRM’s ACT requires a second forward pass for the continue loss, DIS uses a fixed $N_{\text{sup}}{=}6$ with no halting head and no extra forward pass.
@@ 5.2 ARC Evaluation Protocol
Pre-/post-processing. Inputs and outputs are tokenized as discrete color IDs; we concatenate demonstrations and the target input in the same sequence layout used by TRM, our positional scheme match theirs.
For evaluation purposes, we apply the majority vote of the TRM to 1,000 augmented inferences per puzzle. Different number of passes are used and we report pass@2 as it is the most common setting.
@@ 5.3 ARC Model Settings
Objective. Each supervision step $s\in\{1,\dots,6\}$ is trained toward a step-specific intermediate target $\mathbf{y}^{\dagger}_{s}$ produced by a discrete corruption schedule of ground truth $\mathbf{y}^{\star}$ with monotonically decreasing noise. We use token-masking/replacement with a linearly decreasing mask rate over the 6 steps so that $\mathbb{E}[d(\mathbf{y}^{\dagger}_{s},\mathbf{y}^{\star})]$ decreases with $s$. The loss is the standard token-level cross-entropy in $f_{O}(\mathbf{y})$ against $\mathbf{y}_{s}^{\dagger}$, with linearly increasing step weights $w_{s}$ to emphasize late-step fidelity.
@@
Deep improvement supervision loop. For each mini-batch we run $N_{\text{sup}}{=}6$ DIS steps. At each step, we execute a single external cycle (since T=1T{=}1) comprising two internal latent/answer updates (n=2n{=}2), backpropagating through the entire cycle; we then detach $(\mathbf{y},\mathbf{z})$ before the next step. We do not train a halting/ACT head.
Importantly, when using a discrete diffusion model, the supervision trajectories are generated/sampled on the fly, as in a regular diffusion process. Therefore, for the same task, we can have various diffusion steps towards the target.

### v4

@@ 1 Introduction
The goal of reasoning models is to start from a set of specific examples or observations and to infer a general rule or pattern. Such models systematically manipulate and connect pieces of information to infer new conclusions and solve problems not explicitly stated in the training data.
Recently, the use of latent reasoning and iterative refinement loops has become a major driver of progress in deep learning reasoning. Looped transformers (Giannou et al., 2023; Yang et al., 2023) repeatedly apply the same transformer block with input injection at each step and achieve better performance than a standard forward pass transformer in reasoning and meta learning tasks, while using a 10x lower number of parameters (Yang et al., 2023).
@@
TRMs are difficult to train efficiently: supervision is primarily applied to the final output, leaving early refinement steps with weak credit assignment and making optimization brittle in the small-data regime. We propose Deep Improvement Supervision (DIS), a training method that supplies step-wise intermediate targets during recursion. In our main implementation, these targets are generated by a monotonic discrete corruption schedule of the ground-truth output, turning each recursion step into a supervised sub-goal.
@@ 2.1 Hierarchical Reasoning Models
$\mathbf{z}^{t+1}_{L}$ $\leftarrow f^{\phi}_{L}((\mathbf{z}^{t}_{L}+\mathbf{x})+\mathbf{z}^{t}_{H}),\quad\text{(repeated $t$ times)}$ (1)
$\mathbf{z}^{t+1}_{H}$ $\leftarrow f^{\psi}_{H}(\mathbf{z}^{t+1}_{L}+\mathbf{z}^{t}_{H}),$
@@ 2.2 Tiny Recursive Models
The TRM paper explicitly reframes this: $\mathbf{z}_{H}$ is simply the current (embedded) solution… $\mathbf{z}_{L}$ is a latent feature that does not directly correspond to a solution but can be transformed into one by $f_{H}$. It was shown on the Sudoku example that when you reverse-embed + argmax, the tokenized $\mathbf{z}_{H}$ looks like the solved grid, while tokenized $\mathbf{z}_{L}$ is not realted to the solution.
@@ 2.3 Policy Improvement Algorithm
Recently, it was shown that the policy improvement framework does not necessarily rely on explicitly learning a value function. Recent work formalizes a tight connection between classifier-free guidance (CFG) in diffusion/flow models and policy improvement in reinforcement learning (Frans et al., 2025). This work provides a framework for the analysis of diffusion models as RL methods. For a state $\mathbf{s}$ and an action/output $\mathbf{a}$, the classifier-free guidance RL (CFGRL) method parameterizes a target policy as:
@@ 3 Formalization of Latent Reasoning as Policy Improvement
Both TRMs and HRMs are often motivated by effective depth: reusing a small parameter-shared network is said to emulate very deep computation with few parameters. In HRMs this story is tied to Deep Equilibrium (DEQ) models, where one assumes convergence to a fixed point and invokes the Implicit Function Theorem to justify differentiating through only the last iterate (Bai et al., 2019). However, follow-up analyses indicate that the fixed-point assumption is typically violated in practice: latent residuals remain non-negligible under truncation, and a one-step DEQ approximation does not fully explain the empirical gains (Jolicoeur-Martineau, 2025).
@@
$s_{t}\;=\;(x,\mathbf{z}_{L}^{t},\mathbf{z}_{H}^{t}),$
where $\mathbf{x}$ is the embedded input, $\mathbf{z}_{L}^{t}$ is a latent scratchpad, and $\mathbf{z}_{H}^{t}$ is the embedding of the current solution. A parameter-shared backbone is used in two roles:
$\mathbf{z}_{L}^{t+1}=f_{\phi}^{(L)}(\mathbf{x},\mathbf{z}_{H}^{t},\mathbf{z}_{L}^{t}),\qquad\mathbf{z}_{H}^{t+1}=f_{\phi}^{(H)}(\mathbf{z}_{H}^{t},\mathbf{z}_{L}^{t+1}),$ (6)
and an output head $f_{O}$ maps $\mathbf{z}_{H}$ to logits over the next token/action set $\mathcal{A}$. Only $\mathbf{z}_{H}$ is directly decoded and trained with cross-entropy.
@@ 3.1 Two policies per recursion answer
Reference policy (no-reasoning). Forward decoding:
$\ell_{u}^{t}\;:=\;f_{O}(\mathbf{z}_{H}^{t}),\qquad\hat{\pi}_{t}(a\mid s_{t})\;:=\;\mathrm{softmax}(\ell_{u}^{t}).$ (7)
@@
$\ell_{c}^{t}\;:=\;f_{O}(\mathbf{z}_{H}^{t+1}),\qquad\pi_{t}^{+}(a\mid s_{t})\;:=\;\mathrm{softmax}(\ell_{c}^{t}).$ (8)
Intuitively, $\hat{\pi}_{t}$ is the model’s first guess at step $t$, while $\pi_{t}^{+}$ is the refined distribution after one reasoning step. We now introduce the same object used in policy-improvement analyzes: an optimality (or improvement) variable. Let $o\in\{0,1\}$ indicate whether an action $a$ is good [^1] For supervised next-token prediction, the simplest choice is $o=1$ iff $a$ equals the ground-truth token; more general choices can encode task-level reward or self-consistency. for state $s$. A standard KL-regularized improvement step produces a new policy by reweighting a reference policy $\hat{\pi}$ using an optimality likelihood $p(o=1\mid s,a)$:
@@
is an improvement over $\hat{\pi}$. Please, see Remark 1 and Theorem 1 (Frans et al., 2025). We can say that (11) is a special case of (12) with $f(u)=\exp(wu)$.
@@ 3.2 Improvement from a conditional policy
This relationship is evident in Eq. 6, where the updated policy $\pi_{t}^{+}$ depends on $\mathbf{z}_{H}^{t+1}$which is updated via $\mathbf{z}_{H}^{t+1}=f_{\phi}^{(H)}(\mathbf{z}_{H}^{t},\mathbf{z}_{L}^{t+1})$. Crucially, $\mathbf{z}_{L}^{t+1}=f_{\phi}^{(L)}(\mathbf{x},\mathbf{z}_{H}^{t},\mathbf{z}_{L}^{t})$ is explicitly conditioned on the input embedding $\mathbf{x}$. In contrast, standard one-step feed-forward inference does not receive this additional conditioning on $\mathbf{x}$. We can say that the policy after reasoning steps behaves like a policy conditioned on the event that the answer is improved. Under this interpretation, we can recover an optimality by Bayes’ rule:
@@
Equation (16) is the policy-improvement analog of the Bayes inversion step: it rewrites a multiplicative optimality reweighting in terms of a pair of policies (unconditional vs. optimality-conditioned). Concretely, latent reasoning provides both factors $\hat{\pi}_{t}$ and $\pi_{t}^{+}$, and therefore implicitly defines the entire one-step improvement family $\{\pi_{t,w}\}_{w\geq 0}$. From (14), the improvement score is the log-ratio
@@ 3.3 A supervised improvement criterion
For the prediction of the next-token, let $y^{\star}$ be the ground-truth token at the current position. Define the cross-entropy under the improved policy family (16):
@@
Thus, increasing $w$ (trusting the reasoning step more strongly) decreases the loss if and only if the ground-truth action has above-average improvement score:
@@ 4 Deep Improvement Supervision
Section 3 reframes a latent-reasoning step as a policy-improvement operator: each recursion produces a reference policy $\hat{\pi}_{t}(\cdot\mid s_{t})$ and a post-reasoning policy $\pi_{t}^{+}(\cdot\mid s_{t})$, and their log-ratio $A_{t}(s_{t},a)=\log\pi_{t}^{+}(a\mid s_{t})-\log\hat{\pi}_{t}(a\mid s_{t})$ acts as an advantage-like improvement signal. The analysis yielded a concrete Advantage Margin condition: a latent update helps exactly when the ground-truth action has an above-average improvement score under the interpolated policy family $\pi_{t,w}$ (Eq. (18)).
@@
In a recursive architecture, the input to step $s$ (which produces $\ell_{u}^{s}$) is a deterministic function of the post-reasoning state of the previous step (which produces $\ell_{c}^{s-1}$). Under teacher-forced recursion, training $\ell_{c}^{s-1}$ toward $\mathbf{y}^{\dagger}_{s-1}$ implicitly anchors $\ell_{u}^{s}$ near the previous target. Thus, we use the single-loss
@@ 4.1 DIS as direct optimization of the Advantage Margin
Training $\ell_{c}$ on $\mathbf{y}^{\dagger}_{s}$ while $\ell_{u}$ is anchored to $\mathbf{y}^{\dagger}_{s-1}$ forces the residual $\Delta\ell^{s}=\ell_{c}^{s}-\ell_{u}^{s}$ encode the direction of incremental improvement from $\mathbf{y}^{\dagger}_{s-1}$ to $\mathbf{y}^{\dagger}_{s}$. Equivalently, in probability space, DIS shapes the log-ratio advantage $A_{s}(a)=\log\pi_{s}^{+}(a\mid s)-\log\hat{\pi}_{s}(a\mid s)$ assign higher-than-average advantage to the next target token(s), aligning with the theoretical Advantage Margin condition (Eq. (18)). This converts the previous section’s existence/diagnostic statement (“improvement occurs iff the margin is positive”) into a constructive training rule: make the margin positive at every step by supervising the post-reasoning policy on a strictly improving target sequence.
@@ 4.2 Algorithm
Standard TRM training backpropagates a single terminal supervision signal through the entire recursive chain, so the model must simultaneously (i) discover useful intermediate states and (ii) align each latent update with improvement. DIS isolates the second requirement by supervising the improvement operator itself: each recursion step is trained to map the previous target $\mathbf{y}^{\dagger}_{s-1}$ to the next target $\mathbf{y}^{\dagger}_{s}$, directly instantiating the stepwise improvement from § 3.
@@
    Programmable edits. A deterministic generator creates a path from input to output, e.g. puzzle solvers that reveal a constant move in constraints per step, which yields $\mathbf{y}^{\dagger}_{s+1}=\mathrm{Edit}(\mathbf{y}^{\dagger}_{s})$.
@@
    Discrete corruption schedule. Define a corruption process $q_{\beta}(\tilde{\mathbf{y}}\,|\,\mathbf{y}^{\star})$ (e.g., token masking or random replacement with rate $\beta$). Choose a decreasing noise schedule and sample $\mathbf{y}^{\dagger}_{s}\sim q_{\beta_{s}}(\cdot\,|\,\mathbf{y}^{\star})$ so that the targets become progressively less corrupted, yielding a simple monotone-improvement curriculum. [^2] We do not position DIS as diffusion modeling: DIS does not require diffusion schedulers, score matching, or explicit density-ratio estimation (Lou et al., 2023). A corruption schedule is merely a convenient way to generate monotone intermediate targets in discrete spaces.
@@
During training, the model runs the TRM recursion for the $N_{\text{sup}}$ supervision steps. At each step $s$, it computes the post-reasoning logits $\ell_{c}^{s}$ and applies $\mathrm{CE}(\ell_{c}^{s},\mathbf{y}^{\dagger}_{s})$. By the theory in § 3, this stepwise supervision trains each latent update to increase the relative log-probability of the next improved target – i.e., to implement a non-trivial policy improvement operator rather than an unconstrained recurrent computation.
@@ 5 Experiments
From algorithmic perspective, we incorporate the time step $t$ into the model input and observe that an integer-based time index $(0,1\dots,N_{sup})$ yields superior results compared to standard continuous diffusion time conditioning $t\in[0,1])$. The total simplification that we made in comparison to the original TRM is as follows:
@@
  • Halting/ACT: Although TRM/HRM uses halting, where HRM’s ACT requires a second forward pass for the continue loss, DIS uses a fixed $N_{\text{sup}}{=}6$ without halting head and no extra forward pass.
@@ 5.2 ARC Evaluation Protocol
Pre-/post-processing. Inputs and outputs are tokenized as discrete color IDs; we concatenate demonstrations and the target input in the same sequence layout used by TRM, our positional scheme match theirs. For evaluation purposes, we apply the majority vote of the TRM to 1,000 augmented inferences per puzzle. Different number of passes are used and we report pass@2 as it is the most common setting.
@@ 5.3 ARC Model Settings
Objective. Each supervision step $s\in\{1,\dots,6\}$ is trained toward a step-specific intermediate target $\mathbf{y}^{\dagger}_{s}$ produced by a discrete corruption schedule of ground truth $\mathbf{y}^{\star}$ with monotonically decreasing noise. We use token-masking/replacement with a linearly decreasing mask rate over the 6 steps so that $\mathbb{E}[d(\mathbf{y}^{\dagger}_{s},\mathbf{y}^{\star})]$ decreases with $s$. The loss is the standard token-level cross-entropy in $f_{O}(\mathbf{y})$ against $\mathbf{y}_{s}^{\dagger}$.
@@
Deep improvement supervision loop. For each mini-batch we run $N_{\text{sup}}{=}6$ DIS steps. At each step, we execute a single external cycle (since T=1T{=}1) comprising two internal latent/answer updates (n=2n{=}2), backpropagating throughout the cycle; then we detach $(\mathbf{y},\mathbf{z})$ before the next step. We do not train a halting/ACT head. Importantly, when using a discrete diffusion model, the supervision trajectories are generated/sampled on the fly, as in a regular diffusion process. Therefore, for the same task, we can have various diffusion steps towards the target.